OpenStudy (anonymous):

Will give a medal! A line has a slope of -3 and passes through the point (5, -1). What is the equation of the line?

OpenStudy (anonymous):

please help me!

OpenStudy (anonymous):

will give medal and be a fan!

OpenStudy (astrophysics):

You can use the point slope form \[y-y_1=m(x-x_1)\] where your point is \[(x_1,y_1)\] and slope is m, then you can do some algebra once you plug in the values to put it in equation of a line form which is y=mx+b.

OpenStudy (anonymous):

um ok but i need it in equation form not slope intercept form

OpenStudy (anonymous):

is it -3x + y = -2?